# ChatGPT in Data Visualization Education: A Student Perspective

### v1

@@ V-E Conversation Log Analysis Results
We identified three broad themes for user queries: 1) solution-oriented, 2) abstract and open-ended, and 3) interpersonal relational (see Figure 3). The overall theme of each query was determined by the themes of the five codes associated with that query.
A solution-oriented query is one where the user anticipates a specific and detailed response, with a clear goal in mind. The direction is unambiguous, meaning that the system attempt to deliver the sought-after answers or solutions to a posed question. This resembles using ChatGPT as a function where, for a given input X (the query), a corresponding output Y (the response) is produced. The representative codes include: clarification (215/17,835), implementation skills (190), course content (158), implementation (131), customization (131), where the number in the parenthesis represents their frequency. The example queries in this theme include the following: “help me create html to have two boxes so that i can put an image and text in each box”, “i have a csv file with a gender column and a year column I want to add a new column that show the frequency of each gender appears in each years”, “I am attempting to label points on my scatterplot with this code but it isn’t working. What is the problem? $<$CODE$>$”, “using np.isnan replace nan with 0”.
@@
We used the two dominant themes—abstract and open-ended queries, and solution-oriented queries—to categorize the student participants. Our observations revealed that 19 participants more frequently used solution-oriented queries, while eight participants favored abstract and open-ended queries. We analyzed the distribution of themes across different types of assignments. For all assignments, solution-oriented queries outnumbered abstract and open-ended queries. Specifically, the proportion of solution-oriented queries was 61.6% for A1, 64.6% for A2, 51.3% for A3, and 55.1% for A4. Conversely, abstract and open-ended queries comprised 38.4% for A1, 35.4% for A2, 48.7% for A3, and 44.9% for A4. The largest disparity between the two themes occurred in A2 (a 29.3% difference), while the smallest was in A3 (a mere 2.5% difference).
@@ V-F Integrated Analysis Results
We analyzed the correlation between students’ grades and their interactions with ChatGPT, as well as the number of characters in their queries across four assignments. For interactions, A1 showed a slight negative correlation ($R^{2}$=0.008), A2 had no correlation ($R^{2}$=0.000), and A3 and A4 showed slight positive correlations ($R^{2}$=0.023 and $R^{2}$=0.049, respectively). Overall, interactions had minimal impact on grades. In terms of query length, after preprocessing to replace codes, data, and links with placeholders, both A1 and A2 displayed slight negative correlations ($R^{2}$=0.053 and $R^{2}$=0.004, respectively), while A3 and A4 showed slight positive correlations ($R^{2}$=0.013 and $R^{2}$=0.024, respectively).
@@ VI Limitations
Our work has several limitations. The participation was voluntary, possibly introducing selection bias; the participants may have had pre-existing favorable or unfavorable views towards ChatGPT, though they constituted a majority (26 out of 40) of the total enrollment. Additionally, the offer of extra credit for participation could have influenced their motivation. The potential for grade inflation, due to lenient grading practices by teaching assistants, might have masked significant findings in our integrated analysis.
Given that LLMs are continually updated, the study’s results may be specific to the version of ChatGPT available at the time. Some students may have used a paid version of ChatGPT, which includes advanced features such as the ability to read images and datasets. Furthermore, although participants were encouraged to maintain the same chat channel for each assignment, there is a possibility that some engaged in multiple chats (e.g., restarting a session to reset memory), which further complicates the consistency of data collection.
@@ VII Discussion
The unanimous positive reception of ChatGPT highlights its significant potential for LLM-based chatbots in education. Participants appreciated the versatility and efficiency of ChatGPT in assisting primarily with coding tasks, as well as its availability outside of the classroom, reiterating previously reported benefits of chatbots. Moreover, our results point to additional benefits, including ChatGPT’s role in helping students get started and facilitating exploratory learning.
VII-A Design considerations for a tailored education system for data visualization
Several challenges raised by student participants present opportunities for the development of tailored educational solutions, particularly for visually-oriented courses such as data visualization. We discuss various design considerations for these systems below.
@@ VII-A Design considerations for a tailored education system for data visualization
Our study indicates that students rarely inquired about data visualization design knowledge, with only a few engaging in basic chart suggestions. Although not critically reflected in their grades, in-class design feedback sessions revealed that students frequently made less effective design choices, such as using continuous scaling for categorical data and inconsistent color schemes across multiple charts. While the capability of ChatGPT or LLMs in visualization knowledge remains largely unexplored, previous work demonstrates that ChatGPT can provide design feedback on creative work [30, 14]. Although the current LLM-chatbot is constrained by its chat interface, future educational systems could offer user interfaces that better support design-oriented tasks.”
@@
Many frustrations arose from the lack of fluid conversations, often exacerbated by the text-oriented mode of communication, which made it difficult to contextualize participants’ design intentions. Although some students who paid for enhanced services could upload images or datasets, there was still a demand for further improvements in input and output capabilities. For example, participants sought file-reading capabilities, such as processing a set of JavaScript, HTML, and CSS files together and executing the output for visual confirmation. Additionally, the ability to visually annotate subregions in images could further clarify user intents. Supporting such multimodal interaction in future systems can help address the concern expressed by P7, who stated, “I didn’t know the technical terminology or vernacular to be able to fully describe my intentions.”
@@
Many students viewed ChatGPT merely as a simple tool rather than a learning aid as P4 noted that ChatGPT “just gave me the code,” it did not facilitate a deeper understanding of “how it works,” highlighting a gap in fostering learning and creativity. The tool’s effectiveness in providing answers led to concerns about developing an overreliance on it. Future tailored solutions should aim to mitigate this issue by offering progressive learning opportunities that emphasize intermediate steps toward the final solution and promote metacognitive strategies. Additionally, tailoring the degree of assistance based on the student’s level of knowledge can prevent the introduction of incomprehensible advanced concepts or overly complex techniques in the response output.

### v2

@@ V-E Conversation Log Analysis Results
We identified three categories of user queries: 1) solution-oriented, 2) abstract and open-ended, and 3) interpersonal relational (Figure 3). Although we focus on questions asked rather than behavioral characteristics, the first two categories align with Barke et al.’s ‘acceleration’ and ‘exploration’ concepts [7], which echo dual-process theories of cognition [12].
A solution-oriented query expects a specific and detailed response with a clear goal. The direction is unambiguous, prompting the system to deliver the desired solution. This is akin to using ChatGPT as a function, where input X (the query) produces output Y (the response). The representative codes include: clarification (215/17,835), implementation skills (190), course content (158), implementation (131), customization (131), where the number in the parenthesis represents their frequency. The example queries in this theme include the following: “help me create html to have two boxes so that i can put an image and text in each box”, “i have a csv file with a gender column and a year column I want to add a new column that show the frequency of each gender appears in each years”, “I am attempting to label points on my scatterplot with this code but it isn’t working. What is the problem? $<$CODE$>$”, “using np.isnan replace nan with 0”.
@@
We categorized participants based on two dominant query types: abstract/open-ended and solution-oriented. 19 participants primarily used solution-oriented queries, while eight favored abstract/open-ended queries. We analyzed theme distribution across different assignment types. Solution-oriented queries outnumbered abstract/open-ended queries for all assignments. Specifically, the proportion of solution-oriented queries was 61.6% for A1, 64.6% for A2, 51.3% for A3, and 55.1% for A4. Conversely, abstract and open-ended queries comprised 38.4% for A1, 35.4% for A2, 48.7% for A3, and 44.9% for A4. The largest disparity between the two themes occurred in A2 (a 29.3% difference), while the smallest was in A3 (a mere 2.5% difference).
@@ V-F Integrated Analysis Results
We analyzed the correlation between students’ grades and ChatGPT usage (turn-taking interaction frequency and query character count) across four assignments. For interactions, A1 showed a slight negative correlation ($R^{2}$=0.008), A2 had no correlation ($R^{2}$=0.000), and A3 and A4 showed slight positive correlations ($R^{2}$=0.023 and $R^{2}$=0.049, respectively). Overall, interactions had minimal impact on grades. In terms of query length, after preprocessing to replace codes, data, and links with placeholders, both A1 and A2 displayed slight negative correlations ($R^{2}$=0.053 and $R^{2}$=0.004, respectively), while A3 and A4 showed slight positive correlations ($R^{2}$=0.013 and $R^{2}$=0.024, respectively).
@@ VI Limitations
Our work has several limitations: participation was voluntary, possibly introducing selection bias; participants may have had pre-existing favorable or unfavorable views toward ChatGPT; extra credit for participation could have influenced motivation; and the high concentration of grades in the 90th percentile reduces variability, making it challenging to discern differences or achieve statistical significance.
Our study’s results may be specific to the ChatGPT version used at the time. Some students may have used the paid version with advanced features like image and dataset reading. Additionally, despite instructions to use the same chat channel for each assignment, some may have started multiple chats, complicating data consistency.
@@ VII Discussion
The unanimous positive reception of ChatGPT underscores its potential for LLM-based chatbots in education. Participants valued its versatility and efficiency in coding tasks and its availability outside the classroom, echoing benefits reported in related work (Section II) and other professional surveys [37, 51]. Our results also highlight ChatGPT’s benefits in helping students get started and facilitating exploratory learning in a project-oriented data visualization course.
VII-A Considerations for AI-Enhanced Learning Systems
Challenges raised by students highlight opportunities for tailored educational solutions, especially for visually oriented courses like data visualization. We discuss design considerations for these systems below.
@@ VII-A Considerations for AI-Enhanced Learning Systems
Our study indicates that students rarely inquired about data visualization design knowledge, with only a few engaging in basic chart suggestions. Although not critically reflected in their grades, in-class design feedback sessions revealed that students frequently made less effective design choices, such as using continuous scaling for categorical data and inconsistent color schemes across multiple charts. While the capability of ChatGPT or LLMs in visualization knowledge remains largely unexplored, previous work demonstrates that ChatGPT can provide design feedback on creative work [32, 16]. Although the current LLM-chatbot is constrained by its chat interface, future educational systems could offer user interfaces that better support design-oriented tasks.
@@
Many frustrations arose from the lack of fluid conversations, often exacerbated by the text-oriented mode of communication, which made it difficult to contextualize participants’ design intentions. Although some students who paid for enhanced services could upload images or datasets, there was still a demand for further improvements in input and output capabilities. For example, participants sought file-reading capabilities, such as processing a set of JavaScript, HTML, and CSS files together and executing the output for visual confirmation. Additionally, the ability to visually annotate subregions in chart images could further clarify user intents. Supporting such multimodal interaction in future systems can help address the concern expressed by P7, who stated, “I didn’t know the technical terminology or vernacular to be able to fully describe my intentions.”
@@
Many students viewed ChatGPT as a simple tool rather than a learning aid, with P4 noting that it “just gave me the code” without facilitating a deeper understanding of “how it works.” This highlights a gap in fostering learning and creativity, raising concerns about potential overreliance. Future tailored solutions should offer progressive learning opportunities that emphasize intermediate steps and tailor assistance to students’ knowledge levels to overcome incomprehensible responses. Moreover, it would be crucial to define clear boundaries where LLM assistance is beneficial and where students need to develop their own skills, by promoting meta-learning strategies or student-in-the-loop learning [49]. This balance is essential for an effective learning aid, complementing rather than replacing students’ development.
@@ VIII Conclusion
IX Acknowledgement
This research was supported by the National Science Foundation under Grant No. 2146868.